OpenStudy (anonymous):

-t/8=-9 A. t= -72 B. t = 1.125 C. t = 72 D. t = -17

OpenStudy (anonymous):

i think its A

OpenStudy (compassionate):

-t -- = -9 8 Now just multiply: 8*-9 = -72 You're correct.

OpenStudy (anonymous):

yes:) !

OpenStudy (anonymous):

-5a=5/9

OpenStudy (anonymous):

A.1/9 B. -1/9 C.25/9

OpenStudy (anonymous):

D. X=-9

OpenStudy (compassionate):

5 5 - 5 -25 -5a = --- = a = --- x --- ...:::... a = ---- 9 9 1 9

OpenStudy (anonymous):

OKAY SO just multiply

OpenStudy (anonymous):

5/6d=5/9

OpenStudy (compassionate):

Yes, the answer is C

OpenStudy (anonymous):

5/6X5/9

OpenStudy (anonymous):

25/54? = D

OpenStudy (compassionate):

5 --- 6x^5 ---- 9 Is that correct?

OpenStudy (anonymous):

5/6d=2/3

OpenStudy (anonymous):

Answers are : 1. d=4/5 2.d=5/9 3.d=3/2 4.d=7/9

OpenStudy (anonymous):

i mean a.b.c.d

OpenStudy (compassionate):

5 2 --d = -- = (Reverse by multiplying be reciprocal.) 6 3 The answer is d

OpenStudy (anonymous):

thank you i would of never figured that out i would stuck on that ? for ever

OpenStudy (compassionate):

Would you like me to tell you how I came to the answer?

OpenStudy (anonymous):

yes please

OpenStudy (compassionate):

5 2 --d = -- = (Reverse by multiplying be reciprocal.) 6 3 The reciprocal of a fraction is just that fraction turned upsidedown. I know fractions can look nasty in equations, but all you have to do to get rid of them is just multiply that fraction by itself. 5 So in order to get rid of --, we need to multiply the fraction buy its reciprocal. 6 5 6 Algebraically shown, the reciprocal of -- is --, now when you multiply those two they cancel 6 5 Back to the equation, 5 2 6 5 --d = -- ===> -- * --d (We multiplied by its reciprocal canceling the terms.) 6 3 5 6 Now we have isolated, 'd', on one side, all we need to do is multiply the other side by the reciprocal of 5 -- 6 So we have, (Multiply straight across.) 2 6 12 d = -- * -- === --- = 0.75 = 4/5 3 5 15